OpenStudy (anonymous):

On a journey, a cyclist travelled 1 kilometer in x minutes. On as second journey, the cyclist travelled for y hours at the same average speed as on the first journey. Find an expression, in terms of x and y, for the number of kilometers he travelled on the second journey?

OpenStudy (anonymous):

@FoolForMath @robtobey @SmoothMath @AccessDenied @ash2326

OpenStudy (anonymous):

@radar @stormfire1 @KingGeorge

OpenStudy (radar):

For the first journey, d=rt 1km=rx(minutes) so avg speed is 1(km)/x(min) for the second journey, the distance in km expressed in terms of x and y would be: d(km)=(60y)/x

OpenStudy (anonymous):

thanks i got , that...:)

OpenStudy (radar):

Since two different units are used one is minutes and the other hours, you would convert to similar units, I converted to minutes. Others may convert to hours.

OpenStudy (anonymous):

yes i did thaht...thanks for your help :D

OpenStudy (radar):

Maybe storfire1 has further info regarding this problem.

OpenStudy (stormfire1):

I got the same thing: \[\frac{1km}{60x}=\frac{km}{3600y}\]\[\frac {(3600y)(1km)}{60x}=km\]\[d(km)=\frac{60y}{x}\]

OpenStudy (stormfire1):

I went to seconds...which was more work in retrospect :)

OpenStudy (radar):

There it is in a step by step manner, good job stormfire1

OpenStudy (anonymous):

i have a doubt, since they have given minutes we have to convert them to hours, so we have to divide it right so it becomes 1/(1/60)

OpenStudy (stormfire1):

In the final equation it doesn't matter...the relationship is the same.

OpenStudy (stormfire1):

As you can see, radar and I did it converting things two different ways and came up with the same final equation.

OpenStudy (anonymous):

speed on journey 1 s = d/t s = 1/(1/60)= 60x so substitute that in equation 2 avg speed = d/t 60x = d/y d = 60xy? where did i go wrong?

OpenStudy (stormfire1):

\[1min*\frac{60s}{1min}=60s\]

OpenStudy (anonymous):

okay check if i am now s = d/t s = 1/(x/60) = 60/x substitute that is equation 2 60/x = d/y 60y/x = d

OpenStudy (anonymous):

wait\[s= d/t ---> \frac{1}{\frac{x}{60}}\] 1= kilometer x/60 = time since we are dividing it by 60 so it become 60/x

OpenStudy (stormfire1):

Now you're getting ME confused! If you're traveling 1km/minute...then you're traveling (1/60)km per second. Not 1km per second.

OpenStudy (anonymous):

i hr = 60 minutes so i am converting minutes into hours not seconds

OpenStudy (stormfire1):

AHHH...lemme look again then

OpenStudy (stormfire1):

Well, if you're converting minutes to hours it should work out this way: \[s=\frac{d}{t}=\frac{1km}{x~mins}=\frac{60km}{x~hr}\]\[\frac{60km}{x~hr}=\frac{?km}{y~hrs}\]You still end up with \[\frac{(60km)y~hrs}{x~hrs}=\frac{60y}{x}=d(km)\]

OpenStudy (anonymous):

i also ended up with the same answer check and see?

OpenStudy (stormfire1):

then you're right :P

OpenStudy (anonymous):

can u help me with this Given that 198 = 2 * 3^2 * 11 find the smallest integer, k , such that 198k is a perfect square?

OpenStudy (stormfire1):

probably not...I'm much better at physics :P

OpenStudy (stormfire1):

Post it as a separate question and I'm sure one of the math folks will answer it